OpenStudy (anonymous):

2 cars drive 120 miles from A to B. car 1 is 10mph faster and completes the trip 1hr less timer than car 2. Let s be the speed of the slower car. In terms of s, what is the speed of the faster car? In terms of s, how long does the SLOWER car take to complete the journey? In terms of s, how long does the FASTER car take to complete the journey? Express the equation using answers from question 2& 3.

OpenStudy (anonymous):

a) f=s+10

OpenStudy (anonymous):

I got that then b) y=120/(s+10)+1 ?

OpenStudy (anonymous):

Sorry should have bee {120/(s+10)}+1

OpenStudy (anonymous):

then I think c) x={120/(s+10)} ?

OpenStudy (anonymous):

d) I don't really know about

OpenStudy (anonymous):

i think my mind is working slow cuz i just solve the seocnd one and it is right

OpenStudy (anonymous):

It's all good I've been at this question for about 30 min

OpenStudy (anonymous):

about the b) part t= distance/velocity t= 120/s

OpenStudy (anonymous):

oh yeah. So c) would be {120/(s+10)} -1

OpenStudy (anonymous):

t(faster)=distance/(velocity of faster) t= 120/(s+10)